Dry air is passed through a solution containing 10 g of the solute in 90 g of water and then through pure water. The loss in weight of solution is 2.5g and that of pure solvent is 0.05g. The molecular weight of the solute is
1
Expert's answer
2019-05-07T02:11:33-0400

Solution.

"Ps\/Po = 2.5 g\/(2.5g+0.05g)"


Ps - partial pressure of solution


"(Po-Ps)\/Ps = 0.05\/2.5"

"n(H2O) = m(H2O)\/M(H2O)"

n(H2O) = 90/18 = 5.0 mole


"10\/M = n(solution)"

"(Po-Ps)\/Ps = 0.02"

"0.02 = (10\/M)\/(90\/18)"

M = 100 g/mole

Answer:

M = 100 g/mole